OpenStudy (anonymous):

0.10 mol of sodium is reacted completely with water. The resulting aqueous solution formed is reacted with 1.5mol dm^-3 sulphuric acid. What is the volume of sulphuric acid required for the neutralisation? i) (0.10x1000)/1.5 cm^3 ii) (1.5x1000)/0.10 cm^3 iii) (0.10x1000)/(2x1.5) cm^3

OpenStudy (anonymous):

Idk emcrazy! I'm sorry. :( I'm taking chem this year too

OpenStudy (anonymous):

It's ok. I guess I'll have to wait. >_<

OpenStudy (anonymous):

I'll give you a medal while you wait :)

OpenStudy (anonymous):

Haha. Thanks! :)

OpenStudy (anonymous):

is it X(constant) or multiplying sign

OpenStudy (anonymous):

answer is third one...im sure about it

OpenStudy (anonymous):

\[2Na + 2H_{2}O \rightarrow 2NaOH+H_{2}\] (when sodium reacts with water) u get equal no of moles in Na and NaOH =in this case O.1 mole of NaOH obtained the reaction between NaOH and H2SO4 \[2NaOH+H_{2}SO_{4} \rightarrow Na_{2}SO_{4} + H_{2}O\] then u get 0.05 moles of H2SO4 to react with NaOH then apply \[n=\frac{c \times v}{1000}\] \[\frac{0.05 \times 1000}{1.5} \rightarrow \frac{50}{1.5} \times 2 \rightarrow \frac{100}{3}\]

OpenStudy (mos1635):

1.5 mol*dm^-3=1.5 mole/dm^3=1.5 mol/Lt is n;t that so???

OpenStudy (anonymous):

@thushananth01 : I worked out (iii) as well but the book was telling me (ii). So I thought maybe I was wrong. But then the books can be wrong too sometimes. Thanks anyway! @mos1635 : Yes!

OpenStudy (mos1635):

isn't that molarity?

OpenStudy (mos1635):

n=C*V V=n/c V=0.05/1.5 V=1/30 Lt convert to cm^3 V=100/3 cm^3

OpenStudy (mos1635):

@thushananth01 is wright

OpenStudy (mos1635):

@emcrazy14 is wright

OpenStudy (mos1635):

book is wrong

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

@emcrazy14 dont just depend on answers from the book, do ur question, if u hve justified ur answer then ths grt..

OpenStudy (anonymous):

Haha. Ok. :D